OpenStudy (anonymous):

Evaluate the expression using an appropriate double angle formula. cos^2(15) - sin^2(15)

OpenStudy (anonymous):

cos^2(15) - sin^2(15) = cos 2*15 = cos 30 = (sqrt3)/2

OpenStudy (anonymous):

use of the identtiy cos 2x = cos^2x - sin^2x

OpenStudy (anonymous):

How? could you please explain in detail please?

OpenStudy (anonymous):

there is an established formula: cos 2x = cos^2x - sin^2x which is true for all angles x the reverse is also true: cos^2x - sin^2x = cos2x so if you substitute x = 15 you get cos^2(15) - sin^2(15) = cos(2*15) cos(2*15) = cos 30 now you can use a calculator to find the cosine of 30 you get 0.866 which is a valid answer - i used a special right angled triangle

OpenStudy (anonymous):